A commercial lubricant, Dubdeefordy, has K = SN/SM = 13.1 for immiscible solvent M and N. If 5.7 grams of Dubdeefordy is shaken with 115 mL of solvent M and 45 mL of solvent N, what weight of Dubdeefordy will be found in each layer at equilibrium? Show your calculations.
Answer #1

Total mass of Dubdeefordy = 5.7 g

Let mass of Dubdeefordy in solvent M = x

then, mass of Dubdeefordy in solvent N = 5.7 g - x

Concentration of Dubdeefordy in solvent M = (mass of Dubdeefordy in solvent M) / (volume of solvent M)

Concentration of Dubdeefordy in solvent M = (x) / (115 mL)

Concentration of Dubdeefordy in solvent N = (mass of Dubdeefordy in solvent N) / (volume of solvent N)

Concentration of Dubdeefordy in solvent N = (5.7 g - x) / 45 mL

K = SN / SM

K = (Concentration of Dubdeefordy in solvent N) / (Concentration of Dubdeefordy in solvent M)

13.1 = [(5.7 g - x) / 45 mL] / [(x) / (115 mL)]

13.1 * (45 mL / 115 mL) = (5.7 g - x) / x

5.126 = (5.7 g - x) / x

5.126x = 5.7 g - x

5.126x + x = 5.7 g

6.126x = 5.7 g

x = (5.7 g) / 6.126

x = 0.93 g

Mass of Dubdeefordy in solvent M = x

Mass of Dubdeefordy in solvent M = 0.93 g

Mass of Dubdeefordy in solvent N = 5.7 g - x

Mass of Dubdeefordy in solvent N = 5.7 g - 0.93 g

Mass of Dubdeefordy in solvent N = 4.77 g